OpenStudy (anonymous):

Solve the following system of equations. x + y = 14 -x + y = 16 A. (1, 15) B. (1, -15) C. (-1, 15) D. (6, 6)

OpenStudy (anonymous):

x+y-x+y = 2y = 30 y= 15 x + 15 = 14 x = 14-15 = -1 Answer C: X = -1 Y = 15